OpenStudy (anonymous):

If 15(2x^2-y^2)=7xy, then x:y is ?

OpenStudy (anonymous):

\[15(2x^2-y^2)=7xy\]

OpenStudy (hihi67):

henkjan, did you happen to solve the problem i had, cause i got an answer, and i would like to know if we had the same.

OpenStudy (amistre64):

i dunno, but would we solve for x=0, and then x=1; and compare the results?

OpenStudy (king):

2/sqrt15

OpenStudy (anonymous):

The answer given is 5:6

OpenStudy (amistre64):

is x:y x wrt y tho?

OpenStudy (anonymous):

I think yes

OpenStudy (amistre64):

y = 0; x = ? 15(2x^2)=0; x = 0 y = 1, x = ? 15(2x^2-1)=7x 30x^2 -7x -15 = 0 x = 7 +- sqrt(1849) / 60 = 7+-43/60 = 50/60, or -36/60 i get a 5:6 in there but it might be coincident :)

OpenStudy (king):

but thats just trial and error.....lets try to find a proper method

OpenStudy (amistre64):

trial and error IS a proper method.

OpenStudy (amistre64):

its not eloquent and sophist uses what you know to reach what you dont ...

OpenStudy (amistre64):

sophisticated .... im moving from typos to simply not completing words lol

OpenStudy (king):

i know amistre but fr all probs. it might not be so easy so we shuld try and find a nother method

OpenStudy (amistre64):

that i can agree to ;)

OpenStudy (amistre64):

lets develop some thrms

OpenStudy (amistre64):

can we do derivatives with it? 15(2x^2-y^2) = 7xy 15(4x x' - 2y y') = 7x'y + 7xy' 60x x' - 30y = 7x' y + 7x 60x x' - 7x' y = 30y + 7x x'(60x - 7x) = 30y + 7x x' = (30y + 7x)/(60y - 7x) i wonder if that helps

OpenStudy (anonymous):

I think there must be a simpler way..

OpenStudy (amistre64):

simpler? prolly.

OpenStudy (anonymous):

It might be a ninth grade mathematics, Ratio and proportion

OpenStudy (amistre64):

lol. its been like 30 years since I was in the 9th grade :)

OpenStudy (anonymous):

30 x^2 -15x^2 = 7xy divide by y^2 -> 30 x^2/y^2 -15 = 7x/y 30 (x/y)^2-7(x/y) = 15 use abc-formula to solve for x/y

OpenStudy (king):

what formula;btw i did same and got 2/sqrt15

OpenStudy (anonymous):

(-7 +- sqrt(49 - 4*30*-15))/-60 (b+-sqrt(b^2-4ac)/(-2a) wasn't it?

OpenStudy (king):

no sry i got x\y=25,-18

OpenStudy (anonymous):

I got x/y = 50/60 .. so that seems alright... or it is -0.6

OpenStudy (anonymous):

@henkjan I didn't understand your way of solving.

OpenStudy (king):

yeah henkjan sry am sleepy calc. mistake @abdul:quad.equation formula

OpenStudy (anonymous):

I used the 'quadratic equation' to solve ( http://en.wikipedia.org/wiki/Quadratic_equation#Quadratic_formula). We always call it the abc-formula here, as we rewrite an equation into Ax^2+Bx+c =0 and solve it with this formula

OpenStudy (anonymous):

But how did you make it into a quadratic equation?

OpenStudy (anonymous):

You've got: 30x^2 -15y^2 = 7xy Divide by y^2 to turn all the variables into the form (x/y) 30 (x/y)^2 - 15 = 7 (x/y) Or: 30(x/y)^2 - 7(x/y) - 15 = 0 a = 30 b - -7 c = -15

OpenStudy (anonymous):

I understood, Thanks everyone